# Marginal Debiased Network for Fair Visual Recognition

### v1

@@ IV Experiments
IV-A Datasets
BiasedMNIST [48] is a modified MNIST [52] that contains 70K colored images of ten digits. The background color (bias attribute) is highly correlated with each digit category (target attribute). We set the correlation as $\rho\in\{0.999,0997,0.995,0.99\}$ to evaluate the effectiveness of our MDN under different imbalance ratios.
Corrupted CIFAR-10 [49] is generated by corrupting the CIFAR-10 dataset [53] designed for object classification. It consists of 60K images of 10 classes. We set the corruption as a bias attribute, and the object as a target attribute. We vary the ratio of bias-conflicting samples in training dataest and set it to be {0.5%, 1%, 2%, 5%} to simulate dataset bias during training.
@@ IV-A Datasets
TABLE I: Unbiased accuracy on BiasedMNIST dataset with various target-bias correlations. The best accuracy is indicated in bold red and the second best accuracy is indicated in undelined blue.
Corr Vanilla LNL [20] gDRO [54] $\epsilon$-NCE [23] DI [55] EnD [21] MDN (ours)
0.999 11.8$\pm$0.7 18.2$\pm$1.2 22.1$\pm$2.4 33.2$\pm$3.6 15.7$\pm$1.2 59.5$\pm$2.3 74.3$\pm$1.8
0.997 62.5$\pm$2.9 57.2$\pm$2.2 72.7$\pm$1.4 73.9$\pm$0.8 60.5$\pm$2.2 82.7$\pm$0.3 89.8$\pm$1.0
0.995 79.5$\pm$0.1 72.5$\pm$0.9 83.3$\pm$0.2 83.7$\pm$0.4 89.8$\pm$2.0 94.0$\pm$0.6 94.1$\pm$0.5
0.990 90.8$\pm$0.3 86.0$\pm$0.2 92.3$\pm$0.1 91.2$\pm$0.5 96.9$\pm$0.1 94.8$\pm$0.3 95.9$\pm$0.1
@@
TABLE II: Unbiased accuracy on Corrupted CIFAR-10 dataset with varying ratio of bias-conflicting samples. The best accuracy is indicated in bold red and the second best accuracy is indicated in undelined blue.
Ratio (%) Vanilla HEX [56] EnD [21] ReBias [48] MDN (ours)
0.5 23.1$\pm$1.3 13.9$\pm$0.1 22.9$\pm$0.3 22.3$\pm$0.4 30.3$\pm$0.9
1.0 25.8$\pm$0.3 14.8$\pm$0.4 25.5$\pm$0.4 25.7$\pm$0.2 36.3$\pm$0.6
2.0 30.1$\pm$0.7 15.2$\pm$0.5 31.3$\pm$0.4 31.7$\pm$0.4 45.2$\pm$0.6
5.0 39.4$\pm$0.6 16.0$\pm$0.6 40.3$\pm$0.8 43.4$\pm$0.4 57.8$\pm$1.1
IV-B Implementation detail
Experimental setup. All the experiments are implemented with PyTorch. We employ the four-layer CNN with kernel size 7$\times$7 for BiasedMNIST and ResNet-18 [1] pretrained on Imagenet [57] for the remaining datasets as our backbone architecture. We use the Adam optimizer with learning rates of $10^{-3}$ and $10^{-4}$ for {Corrupted CIFAR-10, CelebA, UTK-Face} and {BiasedMNIST}, respectively. The batch size is set to be 128 and 256 for {BiasedMNIST, CelebA, UTK-Face} and {Corrupted CIFAR-10}. We train the networks for 80, 250, 40, and 40 epochs for BiasedMNIST, Corrupted CIFAR-10, CelebA and UTK-Face, respectively, and the learning rates of margin parameters are $5\times 10^{-3}$, $3\times 10^{-2}$, $5\times 10^{-3}$ and $1\times 10^{-2}$. For CelebA, the size of the image is $224\times 224$, and we apply the augmentation of random flip. For UTK-Face, the size of the image is $64\times 64$, and we apply the augmentation of (1) random resized crop and (2) random flip. The balanced meta-validation set is dynamically re-sampling from training set.
IV-C Synthetic experiments
Results on Corrupted CIFAR-10. We also evaluate our model on Corrupted CIFAR-10 and show the unbiased accuracy results in Table II. Due to the challenging scenario, existing methods exhibit only marginal improvements in unbiased accuracy on Corrupted CIFAR-10. For example, ReBias [48] promotes Hilbert-Schmidt independence between the network prediction and all biased predictions, but it only increases the unbiased accuracy from 39.4% to 43.4% at 5% ratio of bias-conflicting samples. Again, our MDN significantly improves the debiasing performances regardless of the bias severities. Benefiting from the added adaptive margins, MDN obtains an average gain of 12.8% compared with the vanilla method.
TABLE III: Results of mitigating gender bias on CelebA dataset. The best accuracy is indicated in bold red and the second best accuracy is indicated in undelined blue. Here Y and B respectively represent target and bias attributes.
@@ IV-C Synthetic experiments
TABLE IV: Results of mitigating age bias on CelebA dataset. The best accuracy is indicated in bold red and the second best accuracy is indicated in undelined blue. Here Y and B respectively represent target and bias attributes.
@@
IV-D Real-world experiments
Results on CelebA. Table III and IV report the results on CelebA dataset to evaluate the effectiveness of our method in mitigating biases with respect to gender and age, respectively. From the results, we have the following observations. First, the vanilla method cannot perform fairly across different samples, which proves the negative effect of spurious correlations. For example, most of the “attractive” images are women in training dataset. As a result, the vanilla model achieves 89.31% on “attractive” women but only obtains 66.24% on “attractive” men (worst-group accuracy). Second, existing debiased methods can improve the model fairness to some extent. DI [55] trains separated target classifiers for each bias class, so as to make the target prediction independent of the bias attribute. ActiveSD [22] performs the second best in fairness, which learns controllable shortcut features for each sample to eliminate bias information from target features. Finally, our proposed MDN outperforms the comparison methods on most tasks. It achieves 78.01%, 71.35% and 6.97 in terms of unbiased accuracy, worst-group accuracy and EOD when mitigating gender bias, and simultaneously obtains 78.43%, 70.93% and 9.58 when mitigating age bias.
TABLE V: Results of mitigating age bias on UTK-Face dataset. The best accuracy is indicated in bold red and the second best accuracy is indicated in undelined blue. Here Y and B respectively represent target and bias attributes.
@@ IV-D Real-world experiments
TABLE VI: Ablation study on UTK-Face dataset. The best accuracy is indicated in bold. Here Y and B respectively represent target and bias attributes.
@@
Results on UTK-Face. Table V shows the debiased performance on UTK-Face dataset. The observations are similar to the ones on CelebA. 1) The vanilla method without fairness constraint usually exhibits low values of fairness metrics. 2) Some previous debiased methods, e.g., ActiveSD [22] and DI [55], improve model fairness compared with the vanilla method; while others perform poorly, e.g., LNL [20] and EnD [21]. We conjecture that this is because more serious dataset bias of UTK-Face makes it difficult for some simple methods to disentangle bias information from target prediction. 3) Our MDN addresses the bias problem from the perspective of margin. During training, we introduce margin penalty and reshape the decision boundary rather than removing bias information to achieve fairness. From the results, we can see that our MDN obtains better debiased performance, especially it improves the worst-group accuracy from 38.69% to 61.21% compared with ActiveSD when mitigating age bias for gender prediction. It also outperforms RWG [59] which re-weights the sampling probability of each example to achieve group-balance, demonstrating the advantage of the modified margin objective. Additionally, MDN achieves 78.33% in terms of unbiased accuracy, which proves that MDN can guarantee the overall model performance on unbiased test criteria when reducing bias.
IV-E Ablation study
@@
Effectiveness of MEL. We also perform an ablation study with MDN w/o MEL in which the margin parameters are updated by meta-optimization guided by the unbiased loss, rather than by MEL. The definition of unbiased loss is similar to that of unbiased accuracy, and it denotes the average (cross-entropy) loss of different groups. Compared with the vanilla method, MDN w/o MEL can improve the performance of the worst group, and thus achieve fairness. However, it is still inferior to MDN which proves the effectiveness of our proposed MEL.
@@ IV-F Visualization
V Conclusion
@@
However, several limitations, including future works, need to be addressed. First, our method requires full knowledge of the bias labels which are unavailable in most practical scenarios. Our further research will delve into investigating a more relaxed condition of MDN that does not require bias labels. Second, to ensure a fair comparison with other methods without introducing additional data, the balanced meta-validation set is constructed by dynamically re-sampling from the training set in MDN. How to get a better and more effective validation set to guide the learning of margin parameters remains to be explored.

### v2

@@ 4 Experiments
4.1 Datasets
BiasedMNIST [33] is a modified MNIST that contains 70K colored images of ten digits. The background color (bias attribute) is highly correlated with each digit category (target attribute). We set the correlation as $\rho\in\{0.999,0997,0.995,0.99\}$ to evaluate the effectiveness of our MDN under different imbalance ratios.
Corrupted CIFAR-10 [34] is generated by corrupting the CIFAR-10 dataset designed for object classification. It consists of 60K images of 10 classes. We set the corruption as a bias attribute, and the object as a target attribute. We vary the ratio of bias-conflicting samples in training dataest and set it to be {0.5%, 1%, 2%, 5%} to simulate dataset bias during training.
@@ 4.1 Datasets
TABLE I: Unbiased accuracy on BiasedMNIST dataset with various target-bias correlations. The best result is indicated in bold red and the second best result is indicated in undelined blue.
Corr Vanilla LNL [12] ReBias [33] gDRO [37] $\epsilon$-NCE [15] LfF [5] DI [38] EnD [13] MDN (ours)
0.999 11.8$\pm$0.7 18.2$\pm$1.2 26.5$\pm$1.4 22.1$\pm$2.4 33.2$\pm$3.6 15.3$\pm$2.9 15.7$\pm$1.2 59.5$\pm$2.3 74.3$\pm$1.8
0.997 62.5$\pm$2.9 57.2$\pm$2.2 65.8$\pm$0.3 72.7$\pm$1.4 73.9$\pm$0.8 63.7$\pm$20.3 60.5$\pm$2.2 82.7$\pm$0.3 89.8$\pm$1.0
0.995 79.5$\pm$0.1 72.5$\pm$0.9 75.4$\pm$1.0 83.3$\pm$0.2 83.7$\pm$0.4 90.3$\pm$1.4 89.8$\pm$2.0 94.0$\pm$0.6 94.1$\pm$0.5
0.990 90.8$\pm$0.3 86.0$\pm$0.2 88.4$\pm$0.6 92.3$\pm$0.1 91.2$\pm$0.5 95.1$\pm$0.1 96.9$\pm$0.1 94.8$\pm$0.3 95.9$\pm$0.1
@@
TABLE II: Unbiased accuracy on Corrupted CIFAR-10 dataset with varying ratio of bias-conflicting samples. The best result is indicated in bold red and the second best result is indicated in undelined blue.
Ratio (%) Vanilla HEX [39] EnD [13] ReBias [33] LfF [5] ECSGA [40] MDN (ours)
0.5 23.1$\pm$1.3 13.9$\pm$0.1 22.9$\pm$0.3 22.3$\pm$0.4 28.6$\pm$1.3 30.9$\pm$0.8 30.3$\pm$0.9
1.0 25.8$\pm$0.3 14.8$\pm$0.4 25.5$\pm$0.4 25.7$\pm$0.2 33.1$\pm$0.8 36.2$\pm$0.4 36.3$\pm$0.6
2.0 30.1$\pm$0.7 15.2$\pm$0.5 31.3$\pm$0.4 31.7$\pm$0.4 39.9$\pm$0.3 44.1$\pm$1.2 45.2$\pm$0.6
5.0 39.4$\pm$0.6 16.0$\pm$0.6 40.3$\pm$0.8 43.4$\pm$0.4 50.3$\pm$1.6 55.1$\pm$1.2 57.8$\pm$1.1
4.2 Implementation detail
Experimental setup. All the experiments are implemented with PyTorch. We employ the four-layer CNN with kernel size 7$\times$7 for BiasedMNIST and ResNet-18 [1] pretrained on ImageNet for the remaining datasets as our backbone architecture. We use the Adam optimizer with learning rates of $10^{-3}$ and $10^{-4}$ for {Corrupted CIFAR-10, CelebA, UTK-Face} and {BiasedMNIST}, respectively. The batch size is set to be 128 and 256 for {BiasedMNIST, CelebA, UTK-Face} and {Corrupted CIFAR-10}. We train the networks for 80, 250, 40, and 40 epochs for BiasedMNIST, Corrupted CIFAR-10, CelebA and UTK-Face, respectively, and the learning rates of margin parameters are $5\times 10^{-3}$, $3\times 10^{-2}$, $5\times 10^{-3}$ and $1\times 10^{-2}$. For CelebA, the size of the image is $224\times 224$, and we apply the augmentation of random flip. For UTK-Face, the size of the image is $64\times 64$, and we apply the augmentation of (1) random resized crop and (2) random flip. The balanced meta-validation set is dynamically re-sampling from the training set. We employ an online re-sampling strategy to ensure that the meta-validation set is continuously updated, which helps prevent the model from overfitting to it to some extent.
TABLE III: Results of mitigating gender bias on CelebA dataset. The best result is indicated in bold red and the second best result is indicated in undelined blue. Here Y and B respectively represent target and bias attributes.
@@ 4.2 Implementation detail
TABLE IV: Results of mitigating age bias on CelebA dataset. The best result is indicated in bold red and the second best result is indicated in undelined blue. Here Y and B respectively represent target and bias attributes.
@@
4.3 Synthetic experiments
Results on Corrupted CIFAR-10. We also evaluate our model on Corrupted CIFAR-10 and show the unbiased accuracy results in Table II. Due to the challenging scenario, existing methods exhibit only marginal improvements in unbiased accuracy on Corrupted CIFAR-10. For example, ReBias [33] promotes Hilbert-Schmidt independence between the network prediction and all biased predictions, but it only increases the unbiased accuracy from 39.4% to 43.4% at 5% ratio of bias-conflicting samples. ECSGA [40] performs the second best in terms of fairness, which reweights bias-aligned and bias-conflicting samples according to their currently produced gradient contributions. Again, our proposed method significantly improves the debiasing performances regardless of the bias severities. Compared with ECSGA [40], MDN obtains an average gain of 0.8% across different ratios, proving the effectiveness of the added adaptive margins.
TABLE V: Results of mitigating age bias on UTK-Face dataset. The best result is indicated in bold red and the second best result is indicated in undelined blue. Here Y and B respectively represent target and bias attributes.
@@ 4.3 Synthetic experiments
4.4 Real-world experiments
Results on CelebA. Tables III and IV report the results on CelebA dataset to evaluate the effectiveness of our method in mitigating biases with respect to gender and age, respectively. From the results, we have the following observations. First, the vanilla method cannot perform fairly across different samples, which proves the negative effect of spurious correlations. For example, most of the “attractive” images in the training dataset are of women. As a result, the vanilla model achieves 89.31% on “attractive” women but only obtains 66.24% on “attractive” men (worst-group accuracy). Second, existing debiased methods can improve the model fairness to some extent. DI [38] trains separated target classifiers for each bias class, so as to make the target prediction independent of the bias attribute. ActiveSD [14], which learns controllable shortcut features for each sample to eliminate bias information from target features, performs best in terms of the EOD metric when using “bag under eye” as the target attribute. Finally, our proposed MDN outperforms the comparison methods on most tasks. It achieves 78.01%, 71.35% and 6.97 in terms of unbiased accuracy, worst-group accuracy and EOD when mitigating gender bias, and simultaneously obtains 78.43%, 70.93% and 9.58 when mitigating age bias.
Results on UTK-Face. Table V shows the debiased performance on UTK-Face dataset. The observations are similar to the ones on CelebA. 1) The vanilla method without fairness constraint usually exhibits low values of fairness metrics. 2) Some previous debiased methods, e.g., ActiveSD [14] and DI [38], improve model fairness compared with the vanilla method; while others perform poorly, e.g., LNL [12] and EnD [13]. We conjecture that this is because more serious dataset bias of UTK-Face makes it difficult for some simple methods to disentangle bias information from target prediction. 3) Our MDN addresses the bias problem from the perspective of margin. During training, we introduce margin penalty and reshape the decision boundary rather than removing bias information to achieve fairness. From the results, we can see that our MDN obtains better debiased performance, especially it improves the worst-group accuracy from 38.69% to 61.21% compared with ActiveSD when mitigating age bias for gender prediction. Compared with the re-sampling methods, MDN outperforms RWG [43] which re-weights the sampling probability of each example to achieve group-balance. Additionally, it successfully reduces EOD from 19.53 to 13.60 compared with BM [41] which is a class-conditioned sampling method. This demonstrates the advantage of the modified margin objective and the effectiveness of considering intra-class variation for bias-conflicting data.
TABLE VI: Ablation study on UTK-Face dataset. The best result is indicated in bold. Here Y and B respectively represent target and bias attributes.
@@ 4.4 Real-world experiments
4.5 Ablation study
@@
Effectiveness of MEL. We also perform an ablation study with MDN w/o MEL in which the margin parameters are updated via meta-optimization guided by the unbiased loss, rather than by MEL. The definition of unbiased loss is similar to that of unbiased accuracy, and it denotes the average (cross-entropy) loss of different groups. Compared with the vanilla method, MDN w/o MEL can improve the performance of the worst group, and thus achieve fairness. However, it is still inferior to MDN which proves the effectiveness of our proposed MEL.
TABLE VII: Conformal inference on BiasedMNIST dataset with various target-bias correlations. The best result is indicated in bold red and the second best result is indicated in undelined blue.
Methods Corr=0.999 Corr=0.997 Corr=0.995 Corr=0.990
Cvg Size CvgD SizeD Cvg Size CvgD SizeD Cvg Size CvgD SizeD Cvg Size CvgD SizeD
($\uparrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\uparrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\uparrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\uparrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$)
Vanilla 0.889 7.845 0.111 2.842 0.903 2.900 0.097 1.753 0.924 1.475 0.076 0.627 0.959 1.322 0.041 0.317
gDRO [37] 0.888 6.810 0.112 4.534 0.916 2.322 0.084 1.385 0.931 1.931 0.069 0.708 0.972 1.241 0.028 0.363
DI [38] 0.894 8.375 0.106 0.864 0.893 2.696 0.082 0.642 0.947 1.071 0.003 0.023 0.978 1.018 0.001 0.001
EnD [13] 0.860 5.613 0.132 0.956 0.883 2.931 0.117 1.778 0.929 2.161 0.071 1.243 0.956 1.484 0.044 0.521
MDN(ours) 0.914 2.138 0.080 1.067 0.961 1.379 0.030 0.416 0.984 1.204 0.014 0.217 0.986 1.224 0.014 0.189
4.6 Analysis
Conformal inference disparity. Conformal prediction [45] is a method of distribution-free uncertainty quantification, which constructs a set of predictions that provably covers the true diagnosis with a high probability. Formally, the uncertainty set function $C(X)$ guarantees states: $P(Y\in C(X))\geq 1-\alpha$, where $X$ and $Y$ are the unseen test example and the true class respectively, and $\alpha$ is a pre-specified confidence level. We conduct conformal inference using RAPS [45] and calculate coverage (Cvg) and set size (Size) on BiasedMNIST dataset to prove the effectiveness of our method. Following [46], we also show coverage disparity (CvgD) and set size disparity (SizeD) to quantify bias. CvgD and SizeD are the differences in Cvg and Size between bias-aligned and bias-conflicting samples. Table VII shows the results on BiasedMNIST dataset, where $\alpha$ is set to be 0.1. As seen from the results, our MDN can obtain satisfactory performance under different target-bias correlations. CvgD and SizeD are successfully reduced to improve fairness, especially when the spurious correlation is higher.
@@ 4.6 Analysis
TABLE VIII: Training cost analysis on BiasedMNIST dataset.
Methods Params GPU memory Training time
Vanilla 0.53 M 1439 MiB ~58 min
EnD [13] 0.53 M 1437 MiB ~123 min
DI [38] 0.54 M 1439 MiB ~59 min
LfF [5] 1.06 M 1641 MiB ~149 min
MDN (ours) 0.53 M 1639 MiB ~97 min
Training cost analysis. Table VIII shows the training cost analysis on BiasedMNIST dataset. Our experiments on BiasedMNIST dataset are implemented in Python on a desktop with one Tesla T4 GPU and Intel Xeon Gold 5218 CPU of 2.3GHz. During training phase, MDN includes both inner loop training and outer loop optimization. The gradient calculations and optimization steps within the inner loop may lead to increased computational overhead compared with the vanilla method. To enhance efficiency, we make $m$ directly only related to the linear classifier rather than the whole classification network, which requires a lightweight overhead. As seen from the results, the increase in computational cost of our proposed MDN is acceptable compared to the baseline model, especially considering the significant performance improvement of MDN. Moreover, the training cost of our MDN is comparable to other existing debiased methods such as LfF and EnD.
5 Conclusion
@@
Limitations and future work. First, our method requires full knowledge of the bias labels which are unavailable in most practical scenarios. Our further research will delve into investigating a more relaxed condition of MDN that does not require bias labels. Second, to ensure a fair comparison with other methods without introducing additional data, the balanced meta-validation set is constructed by dynamically re-sampling from the training set in MDN. How to get a better validation set to guide the learning of margin parameters and avoid overfitting remains to be explored. Third, when the number of bias-conflicting samples in the training set is extremely small, MDN may fail. In the future, we plan to combine our method with generative models.